OpenStudy (1645323):

Find the quadratic equation in standard form with roots -⅔ and 4. Select one: a. 3x2 + 14x + 8 = 0 b. 3x2 -10x - 8 = 0 c. 2x2 - 5x - 12 = 0 d. 2x2 + 9x - 8 = 0

OpenStudy (anonymous):

A

OpenStudy (anonymous):

x = -2/3, x = 4 set to zero combine...

OpenStudy (1645323):

its b @d3Xter

OpenStudy (anonymous):

Oops, sorry about that (sign issues..)

OpenStudy (1645323):

yea

OpenStudy (anonymous):

\[(3x+2)(x-4) \rightarrow 3x^2 -12x+2x -8 = 3x^2-10x-8\]